The average of 6 integers is less than 30. If the average of the first 5 integers is 25, what is the greatest possible value of the 6th integer?
Answer by richwmiller(17219)

5*25=125
The first five integers total 125
If the average of 6 integers is 30 then their sum is 180
180-125=55
The sixth must be 54 or less since we are dealing with integers and 55 makes the average 30
another way to see this is
(5*25+y)/6=30
y=55
the sixth integer must be less than 55

Reminder:
You are limited to one problem per submission.
No more than 4 submissions daily
No similar problems.